# TopoTune: a framework for generalized combinatorial complex neural networks

### v1

@@ 1 Introduction
In this context, Hajij et al. (2023) has recently introduced combinatorial complexes, fairly general objects that are able to model arbitrary higher-order interactions along with a hierarchical organization among them –hence generalizing (for learning purposes) most of the combinatorial topological domains within TDL, including graphs. The elements of a combinatorial complex are cells, being nodes or groups of nodes, which are categorized by ranks. The simplest cell, a single node, has rank zero. Cells of higher ranks define relationships between nodes: rank one cells are edges, rank two cells are faces, and so on. Hajij et al. (2023) also proposes Combinatorial Complex Neural Networks (CCNNs), machine learning architectures that leverage the versatility of combinatorial complexes to naturally model higher-order interactions. For instance, consider the task of predicting the solubility of a molecule from its structure. GNNs model molecules as graphs, thus considering atoms (nodes) and bonds (edges) (Gilmer et al., 2017). By contrast, CCNNs model molecules as combinatorial complexes, hence considering atoms (nodes, i.e. cells of rank zero), bonds (edges, i.e. cells of rank one), and also important higher-order structures such as rings or functional groups (i.e. cells of rank two) (Battiloro et al., 2024).
@@
TopoX (Hajij et al., 2024) and TopoBenchmarkX (Telyatnikov et al., 2024) have become the reference Python libraries for developing and benchmarking TDL models, respectively. However, despite their potential in defining and implementing novel standardized methodologies in the field, the current focus of these packages is on replicating and analyzing existing message-passing CCNNs. Works like Jogl et al. (2022b; a) have tackled this problem by porting TDL models to the graph domain via principled transformations from combinatorial topological domains to graphs. However, although these architectures over the resulting graph-expanded representations are as expressive as their TDL counterparts (using the Weisfeiler-Lehman criterion (Xu et al., 2019a)), the former are neither formally equivalent to nor a generalization of the latter. As such, outside of the expressivity side, the GNNs on the resulting graphs cannot be fairly compared with their TDL counterparts.
@@
  • Implementation. We provide TopoTune, a lightweight PyTorch module for designing and implementing GCCNs. TopoTune is fully integrated into TopoBenchmarkX (Telyatnikov et al., 2024), hosted within the TopoX ecosystem (Hajij et al., 2024). Using TopoTune, both newcomers and expert TDL practitioners can, for the first time, easily define and iterate upon TDL architectures.
  • Benchmarking. Using TopoTune, we create a broad class of GCCNs using four base GNNs and one base Transformer over two combinatorial topological spaces (simplicial and cell complexes). A wide range of experiments on graph-level and node-level benchmark datasets shows GCCNs generally outperform existing CCNNs, often with less model complexity. Some of these results are obtained with GCCNs that cannot be reduced to standard CCNNs, further underlining our methodological contribution. We provide all code and experiment scripts through the TopoBenchmarkX Telyatnikov et al. (2024) repository at github.com/pyt-team/TopoBenchmarkX.
@@ 3 Motivation
The position paper (Veličković, 2022) proposed that any function over a combinatorial complex can be computed via message passing over a transformed graph, but without specifying how to design GNNs that reproduce CCNNs. Later, (Hajij et al., 2023) proposed that, given a combinatorial complex $\mathcal{C}$ and a collection of neighborhoods $\mathcal{N_{C}}$, a message-passing GNN that runs over the augmented Hasse graph $\mathcal{G}_{\mathcal{N_{C}}}$ is equivalent to a specific CCNN as in (3) running over $\mathcal{C}$ using
@@ 5 TopoTune
Our proposed methodology, together with its resulting GCCNs architectures, addresses the challenge of systematically generating principled, general TDL models. Here, we introduce TopoTune, a software module for defining and benchmarking GCCN architectures on the fly—a vehicle for accelerating and democratizing TDL research. TopoTune is made available as part of TopoBenchmarkX Telyatnikov et al. (2024) at github.com/pyt-team/TopoBenchmarkX. This section details TopoTune’s main features.
@@
TopoTune is fully integrated into TopoBenchmarkX (Telyatnikov et al., 2024). This package provides standardized methods for handling combinatorial complex objects, including lifting procedures to lift them from graph-based data. TopoBenchmarkX also simplifies the training and benchmarking of GCCNs, offering a variety of ready-to-use tasks, datasets, and evaluation metrics. Together, TopoTune and TopoBenchmarkX provide all the necessary tools to define, train, test, and compare a wide range of novel GCCN architectures, bringing unprecedented versatility and standardization to accelerate TDL research.
@@ 6 Experiments
We present experiments showcasing a broad class of GCCN’s constructed with TopoTune. These models consistently match, outperform, or improve upon existing CCNNs, often with smaller model sizes. TopoTune’s integration into the TopoBenchmarkX experiment infrastructure ensures a fair comparison with CCNNs from the literature, as implementations of data processing, domain lifting, and training are homogeonized. Moreover, TopoBenchmarkX aggregates TDL’s leading open-source models from TopoX, making it an ideal framework for objectively evaluating GCCNs against the current state of the field.
@@ 6.1 Experimental Setup
We include a wide range of tasks (see Appendix C.2). MUTAG, PROTEINS, NCI01, and NCI09 (Morris et al., 2020) are graph-level classification tasks which seek to predict some property about molecules or proteins. ZINC (Irwin et al., 2012) aims to predict a graph-level property related to molecular solubility through regression. At the node level, the Cora, CiteSeer, and PubMed tasks (Yang et al., 2016) involve classifying publications (nodes) within citation networks. We consider two particular cases of combinatorial complexes, simplicial and cellular complexes, leveraging TopoBenchmarkX’s data lifting processes to infer higher-order relationships in these datasets (see Appendix C.2 for further details).
@@ 6.2 Results and Discussion
Table 1: Cross-domain, cross-task, cross-expansion, and cross-$\omega_{\mathcal{N}}$ comparison of GCCN architectures with top-performing CCNNs benchmarked on TopoBenchmarkX (Telyatnikov et al., 2024). Best result is in bold and results within 1 standard deviation are highlighted blue. Experiments are run with 5 seeds. We report accuracy for classification tasks and MAE for regression.
@@
CCNN (Best Model on TopoBenchmarkX) 80.43 ± 1.78 76.13 ± 2.70 76.67 ± 1.48 75.35 ± 1.50 0.34 ± 0.01 87.44 ± 1.28 75.63 ± 1.58 88.64 ± 0.36
@@
CCNN (Best Model on TopoBenchmarkX) 76.17 ± 6.63 75.27 ± 2.14 76.60 ± 1.75 77.12 ± 1.07 0.36 ± 0.02 82.27 ± 1.34 71.24 ± 1.68 88.72 ± 0.50
@@
CCNN (Best Model on TopoBenchmarkX) 80.43 ± 4.09 76.63 ± 1.74 75.18 ± 1.24 74.93 ± 2.50 0.51 ± 0.01 88.92 ± 0.44 74.93 ± 1.39 89.62 ± 0.25
@@ Appendix B Software
Remark. We decided to design the software module of TopoTune, i.e., how to implement GCCNs, as we did for mainly two reasons: (i) the full compatibility with TopoBenchmarkX (implying consistency of the combinatorial complex instantiations and the benchmarking pipeline), and (ii) the possibility of using GNNs as neighborhood message functions that are not necessarily implemented with a specific library. However, if the practitioner is interested in entirely wrapping the GCCN implementation into Pytorch Geometric or DGL, they can do it by noticing that a GCCN is equivalent to a heterogeneous GNN where the heterogeneous graph the whole augmented Hasse graph, with node types given by the rank of the cell (e.g. 0-cells, 1-cells, and 2-cells) while the edge type is given by the per-rank neighborhood function (e.g. "0-cells to 1-cells" or "2-cells to 1-cells" for $\mathcal{N}^{0}_{I,\uparrow}$ and $\mathcal{N}^{2}_{I,\downarrow}$, respectively).
@@ Appendix D Model Size
CCNN (Best Model on TopoBenchmarkX) 334.72K 101.12K 63.87K 17.67K 88.06K 451.85K 1032.84K 163.72K
@@
CCNN (Best Model on TopoBenchmarkX) 398.85K 10.24K 131.84K 135.75K 617.86K 144.62K 737.29K 134.40K
@@
CCNN (Best Model on TopoBenchmarkX) 84.10K 14.34K 88.19K 88.32K 22.53K 60.26K 258.50K 280.83K

### v2

@@ 1 Introduction
In this context, Hajij et al. (2023; 2024a) have recently introduced combinatorial complexes, fairly general objects that are able to model arbitrary higher-order interactions along with a hierarchical organization among them –hence generalizing (for learning purposes) most of the combinatorial topological domains within TDL, including graphs. The elements of a combinatorial complex are cells, being nodes or groups of nodes, which are categorized by ranks. The simplest cell, a single node, has rank zero. Cells of higher ranks define relationships between nodes: rank one cells are edges, rank two cells are faces, and so on. Hajij et al. (2023) also proposes Combinatorial Complex Neural Networks (CCNNs), machine learning architectures that leverage the versatility of combinatorial complexes to naturally model higher-order interactions. For instance, consider the task of predicting the solubility of a molecule from its structure. GNNs model molecules as graphs, thus considering atoms (nodes) and bonds (edges) (Gilmer et al., 2017). By contrast, CCNNs model molecules as combinatorial complexes, hence considering atoms (nodes, i.e. cells of rank zero), bonds (edges, i.e. cells of rank one), and also important higher-order structures such as rings or functional groups (i.e. cells of rank two) (Battiloro et al., 2024).
@@
TopoX (Hajij et al., 2024b) and TopoBenchmark (Telyatnikov et al., 2024) have become the reference Python libraries for developing and benchmarking TDL models, respectively. However, despite their potential in defining and implementing novel standardized methodologies in the field, the current focus of these packages is on replicating and analyzing existing message-passing CCNNs. Works like Jogl et al. (2022b; a) have tackled this problem by porting TDL models to the graph domain via principled transformations from combinatorial topological domains to graphs. However, although these architectures over the resulting graph-expanded representations are as expressive as their TDL counterparts (using the Weisfeiler-Lehman criterion (Xu et al., 2019a)), the former are neither formally equivalent to nor a generalization of the latter. As such, outside of the expressivity side, the GNNs on the resulting graphs cannot be fairly compared with their TDL counterparts.
@@
  • Implementation. We provide TopoTune, a lightweight PyTorch module for designing and implementing GCCNs. TopoTune is fully integrated into TopoBenchmark (Telyatnikov et al., 2024). Using TopoTune, both newcomers and expert TDL practitioners can, for the first time, easily define and iterate upon TDL architectures.
  • Benchmarking. Using TopoTune, we create a broad class of GCCNs using four base GNNs and one base Transformer over two combinatorial topological spaces (simplicial and cell complexes). A wide range of experiments on graph-level and node-level benchmark datasets shows GCCNs generally outperform existing CCNNs, often with less model complexity. Some of these results are obtained with GCCNs that cannot be reduced to standard CCNNs, further underlining our methodological contribution. We provide all code and experiment scripts through the TopoBenchmark Telyatnikov et al. (2024) repository at github.com/geometric-intelligence/TopoBenchmark.
@@ 3 Motivation
The position paper (Veličković, 2022) proposed that any function over a higher-order domain can be computed via message passing over a transformed graph, but without specifying how to design GNNs that reproduce CCNNs. Later, (Hajij et al., 2023) proposed that, given a combinatorial complex $\mathcal{C}$ and a collection of neighborhoods $\mathcal{N_{C}}$, a message-passing GNN that runs over the augmented Hasse graph $\mathcal{G}_{\mathcal{N_{C}}}$ is equivalent to a specific CCNN as in (3) running over $\mathcal{C}$ using
@@ 5 TopoTune
Our proposed methodology, together with its resulting GCCNs architectures, addresses the challenge of systematically generating principled, general TDL models. Here, we introduce TopoTune, a software module for defining and benchmarking GCCN architectures on the fly—a vehicle for accelerating and democratizing TDL research. TopoTune is made available as part of TopoBenchmark Telyatnikov et al. (2024) at github.com/geometric-intelligence/TopoBenchmark. This section details TopoTune’s main features.
@@
TopoTune is fully integrated into TopoBenchmark (Telyatnikov et al., 2024). This package provides standardized methods for handling combinatorial complex objects, including lifting procedures to lift them from graph-based data. TopoBenchmark also simplifies the training and benchmarking of GCCNs, offering a variety of ready-to-use tasks, datasets, and evaluation metrics. Together, TopoTune and TopoBenchmark provide all the necessary tools to define, train, test, and compare a wide range of novel GCCN architectures, bringing unprecedented versatility and standardization to accelerate TDL research.
@@ 6 Experiments
We present experiments showcasing a broad class of GCCN’s constructed with TopoTune. These models consistently match, outperform, or improve upon existing CCNNs, often with smaller model sizes. TopoTune’s integration into the TopoBenchmark experiment infrastructure ensures a fair comparison with CCNNs from the literature, as implementations of data processing, domain lifting, and training are homogeonized. Moreover, TopoBenchmark aggregates TDL’s leading open-source models from TopoX Hajij et al. (2024b), making it an ideal framework for objectively evaluating GCCNs against the current state of the field.
@@ 6.1 Experimental Setup
We include a wide range of tasks (see Appendix C.2). MUTAG, PROTEINS, NCI01, and NCI09 (Morris et al., 2020) are graph-level classification tasks which seek to predict some property about molecules or proteins. ZINC (Irwin et al., 2012) aims to predict a graph-level property related to molecular solubility through regression. At the node level, the Cora, CiteSeer, and PubMed tasks (Yang et al., 2016) involve classifying publications (nodes) within citation networks. We consider two particular cases of combinatorial complexes, simplicial and cellular complexes, leveraging TopoBenchmark’s data lifting processes to infer higher-order relationships in these datasets (see Appendix C.2 for further details).
@@ 6.2 Results and Discussion
Table 1: Cross-domain, cross-task, cross-expansion, and cross-$\omega_{\mathcal{N}}$ comparison of GCCN architectures with top-performing CCNNs benchmarked on TopoBenchmark (Telyatnikov et al., 2024). Best result is in bold and results within 1 standard deviation are highlighted blue. Experiments are run with 5 seeds. We report accuracy for classification tasks and MAE for regression.
@@
CCNN (Best Model on TopoBenchmark) 80.43 ± 1.78 76.13 ± 2.70 76.67 ± 1.48 75.35 ± 1.50 0.34 ± 0.01 87.44 ± 1.28 75.63 ± 1.58 88.64 ± 0.36
@@
CCNN (Best Model on TopoBenchmark) 76.17 ± 6.63 75.27 ± 2.14 76.60 ± 1.75 77.12 ± 1.07 0.36 ± 0.02 82.27 ± 1.34 71.24 ± 1.68 88.72 ± 0.50
@@
CCNN (Best Model on TopoBenchmark) 80.43 ± 4.09 76.63 ± 1.74 75.18 ± 1.24 74.93 ± 2.50 0.51 ± 0.01 88.92 ± 0.44 74.93 ± 1.39 89.62 ± 0.25
@@ Appendix B Software
Remark. We decided to design the software module of TopoTune, i.e., how to implement GCCNs, as we did for mainly two reasons: (i) the full compatibility with TopoBenchmark (implying consistency of the combinatorial complex instantiations and the benchmarking pipeline), and (ii) the possibility of using GNNs as neighborhood message functions that are not necessarily implemented with a specific library. However, if the practitioner is interested in entirely wrapping the GCCN implementation into Pytorch Geometric or DGL, they can do it by noticing that a GCCN is equivalent to a heterogeneous GNN where the heterogeneous graph the whole augmented Hasse graph, with node types given by the rank of the cell (e.g. 0-cells, 1-cells, and 2-cells) while the edge type is given by the per-rank neighborhood function (e.g. "0-cells to 1-cells" or "2-cells to 1-cells" for $\mathcal{N}^{0}_{I,\uparrow}$ and $\mathcal{N}^{2}_{I,\downarrow}$, respectively).
@@ Appendix D Model Size
CCNN (Best Model on TopoBenchmark) 334.72K 101.12K 63.87K 17.67K 88.06K 451.85K 1032.84K 163.72K
@@
CCNN (Best Model on TopoBenchmark) 398.85K 10.24K 131.84K 135.75K 617.86K 144.62K 737.29K 134.40K
@@
CCNN (Best Model on TopoBenchmark) 84.10K 14.34K 88.19K 88.32K 22.53K 60.26K 258.50K 280.83K